OpenStudy (naveenbhatia1312):

Use basic identities to simplify the expression. (6 points) sin^2θ+ tan^2θ+ cos^2θ

OpenStudy (danjs):

remember the very first one sin^2 + cos^2 = 1 from the unit circle pythagorean theorem

OpenStudy (naveenbhatia1312):

yes

OpenStudy (danjs):

and there is another pythagorean theorem on the unit circle, tan^2 + 1^2 = sec^2

OpenStudy (naveenbhatia1312):

after what you do?

OpenStudy (danjs):

sin^2θ+ tan^2θ+ cos^2θ they just want you to simplify that using the identities, notice it already has sin^2 + cos^2, you can replace that with 1

OpenStudy (danjs):

then use the other pythagorean one i mentioned to finish it up

OpenStudy (naveenbhatia1312):

so it would equal tan theta

OpenStudy (danjs):

you start with sin^2(a) + cos^2(a) + tan^2(a) use the sin^2(a) + cos^2(a) = 1 identity gives 1 + tan^2(a) use the tan^2(a) + 1 = sec^2(a) identity gives sec^2(a)

OpenStudy (danjs):

here is the geometry of where the two identity come from, the pythagorean theorem of these lines